OpenStudy (anonymous):

MEDAL AND FAN!!!! A house cost $120,000 when it was purchased. The value of the house increases by 10% each year. Find the rate of growth each month and select the correct answer below.

OpenStudy (anonymous):

Answer choices 0.797% 0.00797% 0.833% 0.00833%

OpenStudy (itsmichelle29):

D!!!

OpenStudy (itsmichelle29):

LOL its the right answer trust me

OpenStudy (itsmichelle29):

you take 0.10 and divide by 12 to get you answer D.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Solve 5^(2x) = 7^(x−1)

OpenStudy (itsmichelle29):

x=−1.528643

OpenStudy (anonymous):

@fugazi go to ur question

OpenStudy (anonymous):

that is wrong

OpenStudy (itsmichelle29):

HOw

OpenStudy (anonymous):

i got it wrong

OpenStudy (anonymous):

try C.

OpenStudy (anonymous):

that is wrong too

OpenStudy (anonymous):

i think it is B

OpenStudy (anonymous):

yeah b my bad

OpenStudy (anonymous):

the correct answer is B

OpenStudy (anonymous):

You decide to put $100 in a savings account to save for a $3,000 down payment on a new car. If the account has an interest rate of 2% per year and is compounded monthly, how long does it take you to earn $3,000 without depositing any additional funds

OpenStudy (anonymous):

is there any answer choices?

OpenStudy (anonymous):

answer choices 170.202 14.3129 171.755 168.354

OpenStudy (anonymous):

i know c and b is wrong i think it is d

OpenStudy (anonymous):

well to check yourself you can use the formula FV = PV*(1+r/n)^(n*t) the variables are FV = future value PV = present value r = interest rate n = compounding frequency t = time in years You will plug in the following FV = 3000 is the amount you want PV = 100 is the amount you have r = 0.02 n = 12 t = unknown. You are solving for this

OpenStudy (anonymous):

72000

OpenStudy (anonymous):

just plug in the choices for x in 100(1.02)^x

OpenStudy (anonymous):

till u get 3000

OpenStudy (anonymous):

i think its c

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

A motorcycle cost $12,000 when it was purchased. The value of a motorcycle decreases by 6% each year. Find the rate of decay each month and select the correct answer below.

OpenStudy (anonymous):

i got -0.005% is that an answer choice?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

c was wrong a couple of comments ago

OpenStudy (anonymous):

then try that :)

OpenStudy (anonymous):

ur smart